OpenStudy (dbzfan836):

Find -3a^2 - b^3 + 3c^2 - 2b^3 if a = 2, b = -1, and c = 3.

OpenStudy (dbzfan836):

(╯°□°)╯︵ ┻━┻ ┻━┻ ︵ヽ(`Д´)ノ︵ ┻━┻ ┬──┬ ノ( ゜-゜ノ) (ノಠ益ಠ)ノ彡┻━┻

OpenStudy (dbzfan836):

@annas

OpenStudy (anonymous):

coming just a sec

OpenStudy (anonymous):

-12+1+27+2 = -11 + 29 = 29-11 = 18